Grade 12th passAlgebra

The root of z^3+8i=?.......................................

Alina Bhattacharya
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student

z^3 + 8 i = 0
z^3 + (2 i)^3  = 0
(z + 2i)  (z^2 + 2 i^2 – 2*z*i) = 0
I hope now you can solve it on your own.
 
Regards
Arun (askIITians forum expert)